# Understanding LLMs’ Cross-Lingual Context Retrieval: How Good It Is And Where It Comes From

### v1

@@ 4.2.2 Estimated oracle performance
As described in §2.1.3, the oracle performances of the LLaMA models estimated with one generation step and the whole sequence666”Step” means the attribution calculation is done on the probability distribution of the first generation step (generating the first answer token), which roughly measures the ability to locate the correct context; while ”Sequence” means the calculation is done on the multiplied probability distribution of whole-sequence generation (until EOS), which is more realistic. are shown in Table 3, where two phenomena stand out:
First, the estimated oracle performances of post-trained models are significantly higher than the base models, both on en-en and en-x settings, suggesting post-training is crucial to the xMRC task. On the other hand, the advantage of 70B models over 8B models is not as substantial, indicating extended pretraining is expected to contribute less to xMRC ability.
Second, the estimated oracle performance for en-x is close to en-en for all the LLaMA models, and the oracle for post-trained models are basically over 90%. This is quite different from the actual performance, where en-en is better than en-x, and the performances are not that high. This suggests the models have actually obtained the ability to locate the correct answer in the xMRC task, but the ability needs to be further elicited.
@@
5 Two-phased mechanism of xMRC
The above results show that post-training on smaller LLM elicits the cross-lingual transfer of contextual retrieval ability. However, the mechanism that enables such transfer is unknown. Considering the model forward process from the input prompt to the output answer, we come up with a two-phase hypothesis of the xMRC process (taking en-x as an example):
  1. 1.
    Question encoding. The queries (mainly the non-English question) will be encoded into a shared semantic space, where queries in different languages are aligned and understood in a language-neutral way;
  2. 2.
Figure 2 shows an illustration of the hypothesized process. We test our hypothesis from attribution and hidden-state views. Since LLaMA-3.1-Instruct-8B shows high performance alignment across languages and is widely used in the community, we take its family to perform our analysis.
Figure 2: Illustration of the hypothesized two-phased xMRC process. As layer depths increases, the last question token will be gradually transferred to the first answer token.
To further ensure the effect of post-training, we also conduct finetuning and compare the model behavior before and after it.
5.1 Attribution view
We use layer-wise attribution (§2.2.1) to identify the question encoding and answer retrieval phases on the LLaMA models. Figure 3 shows the mean MRD of the contexts and the questions for the LLaMA-3.1-Instruct-8B on testing samples that are identified as either "balanced" or "en-superior" in all tested languages.  [^7] We identify a sample as ”balanced” if the model F1 score on it is above 0.5 in all directions; and a sample as ”en-supeiror” if the F1 score in English is higher than the average of other languages with a margin greater than 0.5.
@@ 5.1 Attribution view
We find that the mean question MRD is significantly and substantially lower than the mean context MRD in all tested languages and across the LLaMA models, especially for the "balanced" samples (Figure 3 and Appendix B.1.2), revealing a clear phased behavior.
Also, by comparing the MRDs of "balanced" (Figures 3(a)) and "en-superior" samples (Figure 3(b)), we can see that the MRDs of "balanced" samples are more stable than those of "en-superior" samples, suggesting correlation between the high cross-lingual context retrieval ability and a clear two-phase behavior.
Details and More results can be found in Appendix B.1. We found this pattern consistent for different LLaMA models and not affected by the prompt format, though it is weaker in LLaMA-2-Chat-7B, which has a smaller pre-trained capacity and weaker multilingual ability.
@@
(a) balanced samples
(b) en-superior samples
Figure 4: Question hidden state similarity between English and other languages in each layer of the LLaMA-3.1-Instruct-8B model.
(a) balanced samples
(b) en-superior samples
Figure 5: Last-input-token hidden state similarity between English and other languages in each layer of the LLaMA-3.1-Instruct-8B model.
(a) balanced samples
(b) en-superior samples
Figure 6: Context hidden state similarity between English and other languages in each layer of the LLaMA-3.1-Instruct-8B model.
5.2 Hidden states view
The hidden state similarity results also support our hypothesis. Figures 4, 5, and 6 show the en-x hidden state similarity of the question, the last input token (predicting the start of the answer) and the context [^8] Although the context parts are all English,there are 2-shot demonstrations in each language direction, making the curves not identical to each other. Since the similarity score refers to cross-language similarity over in-language, the context curves can represent the degree of semantic encoding compared with formatic encoding. parts for the LLaMA-3.1-Instruct-8B model. Results for other LLaMA models can be found in Appendix B.2. The observed trends are consistent:
@@ 5.2 Hidden states view
  • For context representations, a consistent double-peak trend can be observed, with a "turning point" around the relative depth of 0.4 (matching the question MRD) and the second, higher peak at around 0.7 (matching the context MRD);
  • For last input token representations, one can see a consistent "plateau" of similarity starting at around a relative depth of 0.5, which also matches the mean question MRD.
Again, for the less powerful model LLaMA-2-Chat-7B (Figure 22 in Appendix B.2), the trends are weaker: its question similarity to English varies much across languages, and the "plateau" of answer similarity to English starts later than other models, which is after the mean question and context MRDs.
@@
Figure 7: Change in last-input-token hidden state similarity between English and other languages in each layer of LLaMA-3.1-8B, LLaMA-3.1-Instruct-8B and LLaMA-3.1-Tuned-8B on the "balanced" samples.
@@
Figure 8: Change in context hidden state similarity between English and other languages in each layer of LLaMA-3.1-8B, LLaMA-3.1-Instruct-8B and LLaMA-3.1-Tuned-8B on the "balanced" samples.
5.3 Pre-training vs. post-training
Although the phasing behavior is shaped during pre-training, our evaluation results show that post-training is more important to enhance xMRC performance. Here, we show from the hidden-state similarity view the importance of post-training on this task.
The first evidence comes from the last-input-token similarity results. One can observe from Figure 7(a) and 9(a) that, the last-input-token similarity of base models experience severe and consistent decline in the last few layers, across all non-English languages. Since we expect the same English output for all tested languages, this drop in cross-lingual similarity can directly affect the performance and its cross-lingual alignment. However, after post-training (Figure 7(b),7(c) and 9(b)), the decline significantly narrows, and even turns into increase for languages with Latino alphabets, especially on the 8B model (Figure 7). Since this enhancement in similarity can directly turn into the narrowing of language performance gap, this indicates that post-training is especially essential to enhancing the cross-lingual alignment xMRC ability, by taking effect in the last few layers and the final calculation steps. However, for the 70B model, the decline is still severe after post-training, partly explaining why they show larger xMRC gap between English and non-English.
Another evidence comes from the context similarity results. One can see from Figure 6 and 10 that, the models show divergence in context similarity after post-training. Languages more similar to English (e.g. ro, es) tend to diverge lower, and those less similar (e.g. ar, vi) tend to diverge higher. This means demonstrations in languages that are more similar to English will drive the context representation more away from English, possibly making it more suitable for cross-lingual retrieval. Again, this phenomenon is more evident in the 70B model (Figure 10), suggesting its responses to demonstrations in different languages are more diverged, possibly leading to the larger xMRC gap between languages.
In this regard, a possible reason for the larger language gap for 70B post-trained models could be, their post-training is insufficient to reshape the answer retrieving phase, and is biased to English-similar languages. With more sufficient and language-balanced post-training, we expect the Instruct-70B model to reveal similar patterns in last-input-token and context hidden states similarity as Instruct-8B.
@@
Figure 9: Change in last-input-toke hidden state similarity between English and other languages in each layer of LLaMA-3.1-70B and LLaMA-3.1-Instruct-70B on the "balanced" samples.
@@
Figure 10: Change in context hidden state similarity between English and other languages in each layer of LLaMA-3.1-70B and LLaMA-3.1-Instruct-70B on the "balanced" samples.
@@ 6.1 Cross-lingual alignment of LLMs
Previous studies have shown a misalignment of LLMs with English and other languages. With respect to performance, Lai et al. (2023b), Ahuja et al. (2024) and Wang et al. (2024) demonstrated that SOTA LLMs performed better in English, and showed inconsistency when dealing with non-English queries. Etxaniz et al. (2024) found that LLMs performed worse with non-English prompts than with self-translated English prompts. Beyond performance, Qi et al. (2023) demonstrated the low cross-lingual consistency of factual knowledge does of LLMs, and Gao et al. (2024) showed that multilingual pre-training and instruction tuning could only enhance superficial levels of cross-lingual alignment. However, the previous work mostly focused on queries with a consistent language, instead of cross-lingual queries.
There have also been many techniques to enhance LLMs’ cross-lingual alignment. For example, adding parallel data in the pre-training stage Lample and Conneau (2019); Jiang et al. (2022); Wei et al. (2023); Lu et al. (2024); and the post-training stage, including instruction tuning Li et al. (2023b); Zhang et al. (2025); Li et al. (2023a); Cahyawijaya et al. (2023); Chai et al. (2024); Kuulmets et al. (2024); Shaham et al. (2024); Kew et al. (2024) and preference tuning Lai et al. (2023a); She et al. (2024). Especially, extra translation training is commonly used Zhang et al. (2023); Yang et al. (2023b); Li et al. (2024); Ranaldi et al. (2024); Zhu et al. (2024); Lu et al. (2024). In this paper, we examine the effect of some of these techniques by comparing various SOTA models.
@@ 6.2 Cross-lingual machine reading comprehension
xMRC is a relatively new task of natural language understanding. Cui et al. (2019) proposed the task, in order to improve non-English MRC performance by introducing English resources. There are some representative datasets in this area, such as XQA Liu et al. (2019), BiPaR Jing et al. (2019), MLQA Lewis et al. (2020), and XQuAD Artetxe et al. (2020). Ushio et al. (2023) also proposes a pipeline for multilingual QA generation.
Previous work on the xMRC task mainly focuses on enhancing the performance of task-specific models using techniques such as data augmentation Bornea et al. (2021); Xiang et al. (2024), knowledge injection Duan et al. (2021), constractive learning Chen et al. (2022) and knowledge transfer Cao et al. (2023); Xu et al. (2023). However, the xMRC performance of LLMs has not been studied.
@@ 7 Conclusion
We investigate the cross-lingual context retrieval ability of LLMs with the xMRC task. We first evaluates the xMRC performance of existing open- and closed-sourced LLMs, and find that post-trained 7-9B models show high performance and little gap across English and non-English languages. Then, we conduct analysis on the LLaMA models and identifies the two main phases (question encoding and answer retrieval) of the xMRC process, and find a possible explanation to the language gap for larger post-trained models. We hope our research will inspire future study to foster the cross-lingual alignment of LLMs in a broader scope.
@@
While this study provides valuable insights into the cross-lingual context retrieval abilities of LLMs and identifies a two-phased mechanism underlying this process, it is important to acknowledge certain limitations.
First, the scope of our empirical evaluation is constrained by available resources and time. This necessarily limits the breadth of our testing, preventing us from exhaustively covering the rapidly expanding landscape of LLMs. Furthermore, while we test across 12 diverse languages, a more comprehensive analysis would ideally include an even wider range of languages in order to ensure the generalizability of our findings across linguistic diversity.
Second, although we successfully identify a two-phased feature of cross-lingual machine reading comprehension and confirm its correlation with pre-training and post-training stages, the precise factors within these training processes that drive this outcome remain unclear. Future work could delve deeper into the specifics of pre-training objectives, data composition, and post-training techniques to pinpoint the exact elements that contribute to the emergence and effectiveness of these two phases.
Finally, within the two major phases we discover – question encoding and answer retrieval – we observe hints of more fine-grained changes in model behavior, particularly in the hidden state similarity curves. These preliminary observations suggest the potential for a more nuanced understanding of the xMRC process. Future studies could further investigate these finer-grained dynamics within each phase to gain a more detailed and complete picture of how LLMs achieve cross-lingual context retrieval.
@@
Our use of these existing artifacts, including both language models and datasets, is aligned with their intended use within research contexts. Specifically, derivatives of data accessed for research purposes, such as model outputs and analysis results, are used solely within the bounds of academic inquiry and are not disseminated or utilized outside of these research contexts, in accordance with responsible data handling practices. We have striven to conduct this research responsibly and ethically, focusing on understanding and improving the cross-lingual capabilities of language models for the benefit of the NLP community, while respecting the intended use and access conditions of all resources utilized.
@@ Acknowledgments
We would like to thank the anonymous reviewers for their insightful comments. Shujian Huang is the corresponding author. This work is supported by National Science Foundation of China (No. 62376116, 62176120), research project of Nanjing University-China Mobile Joint Institute, the Fundamental Research Funds for the Central Universities (No. 2024300507).
Appendix A Evaluation
A.1 Full list of models evaluated

### v2

@@ 4.2.2 Estimated oracle performance
Table 2 shows the estimated oracle performances of the LLaMA-3.1 models. “Step” means the attribution is done when generating the first answer token, and “Sequence” means the attribution is done on whole-sequence generation until EOS.
First, the estimated oracle performances of post-trained models are significantly higher than the base models, both for en-en and en-x, suggesting the importance of post-training to improving xMRC. However, 70B models show no substantial advantage over 8B, indicating extended pretraining contributes less to xMRC.
Second, the estimated oracle performance for en-x is close to en-en for all the LLaMA-3.1 models, and the oracle for post-trained models are basically over 90%. This is much higher than the actual performance, suggesting the models have the potential to locate correct answers with high accuracy, but the ability needs to be further elicited.
@@
Figure 2: Illustration of the hypothesized two-phased xMRC process. Through layers, the last question token will be transferred to the first answer token in two phases, between which is a cross-lingual question representation.
@@
5 Two-phased mechanism of xMRC
Besides evaluation of xMRC performance and bottlenecks, the mechanism of how LLMs do xMRC also worth exploring. Considering the model forward process from the input prompt to the output answer, we come up with a two-phase hypothesis of the xMRC process (taking en-x as an example):
  1. 1.
    Question encoding. The non-English queries will be encoded into a shared semantic space, where queries in different languages are aligned and understood in a language-neutral way;
  2. 2.
Figure 2 shows an illustration of the hypothesized process, which aligns with previous studies Tang et al. (2024); Wendler et al. (2024); Zhao et al. (2024). If the hypothesis holds, then we will be able to extract cross-lingual representations in the middle layers and steer them to control the model retrieval behaviors. We test our hypothesis from attribution and hidden-state views with LLaMA-3.1-Instruct-8B, which shows high performance alignment across languages and is widely used. To further ensure the effect of post-training, we also conduct finetuning and compare the model behavior before and after it.
5.1 Evidence from attribution view
Figure 3 shows the mean MRD (§2.2.1) of the contexts and the questions for the LLaMA-3.1-Instruct-8B on testing samples that are identified as either “balanced” or “en-superior” in all tested languages. We identify a sample as “balanced” if the model F1 score on it is above 0.5 in all directions; and a sample as “en-superior” if the F1 score in English is higher than the average of other languages with a margin greater than 0.5.
We find that the mean question MRD is significantly and substantially lower than the mean context MRD in all tested languages and across the LLaMA models, especially for the “balanced” samples (Figure 3), revealing a clear phased behavior.
(a) Question
(b) Last token
(c) Context
Figure 4: Question, last token and context hidden state similarity between English and other languages in each layer of the LLaMA-3.1-Instruct-8B model on the “balanced” samples.
Also, the MRDs of “balanced” samples (Figures 3(a)) are more stable than those of “en-superior” samples (Figure 3(b)), suggesting correlation between the higher xMRC performance and clearer phasing.
These patterns are consistent for different LLaMA models regardless of prompt formats, except much weaker for LLaMA-2-Chat-7B, which has smaller pre-training capacity and weaker multilingual ability (see Appendix C.1 for mode details).
@@ 5.1 Evidence from attribution view
5.2 Evidence from hidden states view
The hidden state similarity results also support our hypothesis. Figure 4 shows the en-x hidden state similarity of the question, last input token (predicting the start of the answer) and context parts for the LLaMA-3.1-Instruct-8B model (more results are in Appendix C.2). The observed trends are consistent:
@@ 5.2 Evidence from hidden states view
  • For context representations, a consistent double-peak trend can be observed, with a “turning point” around the relative depth of 0.4 (matching the question MRD) and the second, higher peak at around 0.7 (matching the context MRD);
  • For last input token representations, one can see a consistent “plateau” of similarity starting at around a relative depth of 0.5, which also matches the mean question MRD.
It is worth noticing that, though the context parts are all English, there are 2-shot demonstrations with non-English questions, making the representations not identical. The context similarity curves can represent the degree of semantic encoding compared with formal encoding.
Again, for the less powerful model LLaMA-2-Chat-7B (Figure 20 in Appendix C.2), the trends are weaker: its question similarity to English varies much across languages, and the “plateau” of answer similarity to English starts later than other models, which is after the mean question and context MRDs.
@@
Figure 5: Change in last-input-token hidden state similarity between English and other languages in each layer of LLaMA-3.1-8B, LLaMA-3.1-Instruct-8B and LLaMA-3.1-Tuned-8B on the “balanced” samples.
@@
Figure 6: Change in context hidden state similarity between English and other languages in each layer of LLaMA-3.1-8B, LLaMA-3.1-Instruct-8B and LLaMA-3.1-Tuned-8B on the “balanced” samples.
5.3 Importance of post-training
Our evaluation results show that post-training is crucial to enhancing xMRC performance. Here, we show the significance of post-training to xMRC from the hidden-state view.
The first evidence comes from the last-input-token similarity results. One can observe from Figure 5(a) and 7(a) that, the last-input-token similarity of base models experience severe and consistent decline in the last few layers, across all non-English languages. Since we expect the same English output for all tested languages, this drop in cross-lingual similarity can directly affect the performance and its cross-lingual alignment. However, after post-training on 8B models (Figure 5(b), 5(c) and 7(b)), the decline significantly narrows, and even turns into increase for languages with Latino alphabets, especially on the 8B models (Figure 5). Since this enhancement in similarity can directly turn into the narrowing of language performance gap, this indicates that post-training is especially essential to enhancing the cross-lingual alignment xMRC ability, by taking effect in the last few layers and the final calculation steps. However, for the 70B model (Figure 7), the decline is still severe after post-training, partly explaining why they show larger xMRC gap between English and non-English.
Another evidence comes from the context similarity results. One can see from Figure 6 and 8 that, the context similarity value significantly decreases after post-training, and the peak value diverges among languages. Based on our understanding of the score, lower context similarity means its contextual representations become more customized for the demonstrations with non-English questions, which is potentially beneficial for xMRC. This also partly explains the enhancement of xMRC after post-training. However, the divergence among languages means that this driving effect of demonstrations varies in different languages, contributing to the performance gap between languages. It is especially evident for the 70B model (Figure 8) that, better-xMRC-performing languages (e.g. de, es) tend to show lower peak values, while worse-xMRC-performing languages (e.g. ar, th) tend to show higher. This corresponds with the reasoning that lower context similarity correlates with higher xMRC performance, and adds to the explanation why 70B models have larger performance gaps between languages.
Based on these findings, we propose a possible reason for the larger language gap for 70B post-trained models that their post-training is insufficient to reshape the answer retrieving phase, and is biased to certain languages, which causes the near-base behavior in last-token similarity and the larger divergence in context similarity. With more sufficient and language-balanced post-training, we expect the Instruct-70B model to reveal similar patterns in last-input-token and context hidden states similarity as Instruct-8B.
@@ 5.3 Importance of post-training
Figure 7: Change in last-input-toke hidden state similarity between English and other languages in each layer of LLaMA-3.1-70B and LLaMA-3.1-Instruct-70B on the “balanced” samples.
@@
Figure 8: Change in context hidden state similarity between English and other languages in each layer of LLaMA-3.1-70B and LLaMA-3.1-Instruct-70B on the “balanced” samples.
@@ 6.1 Cross-lingual alignment of LLMs
Previous studies have shown a misalignment of LLMs with English and other languages. With respect to performance, Lai et al. (2023b), Ahuja et al. (2024) and Wang et al. (2024) demonstrated that SOTA LLMs performed better in English, and showed inconsistency when dealing with non-English queries. Etxaniz et al. (2024) found that LLMs performed worse with non-English prompts than with self-translated English prompts. Beyond performance, Qi et al. (2023) demonstrated the low cross-lingual consistency of factual knowledge does of LLMs, and Gao et al. (2024) showed that multilingual pre-training and instruction tuning could only enhance superficial levels of cross-lingual alignment. Zhao et al. (2024) demonstrated that multilingual LLMs employ shared and language-specific circuits to process different languages when needed. These papers mostly focus on same-language queries, and our work focuses on cross-lingual queries.
There have also been many techniques to enhance LLMs’ cross-lingual alignment. For example, adding parallel data in the pre-training stage Lample and Conneau (2019); Jiang et al. (2022); Wei et al. (2023); Lu et al. (2024); and the post-training stage, including instruction tuning Li et al. (2023b); Zhang et al. (2025); Li et al. (2023a); Cahyawijaya et al. (2023); Chai et al. (2024); Kuulmets et al. (2024); Shaham et al. (2024); Kew et al. (2024) and preference tuning Lai et al. (2023a); She et al. (2024). Especially, extra translation training is commonly used Zhang et al. (2023); Yang et al. (2023); Li et al. (2024); Ranaldi et al. (2024); Zhu et al. (2024); Lu et al. (2024). In this paper, we examine the effect of some of these techniques by comparing various SOTA models.
@@ 6.2 Cross-lingual machine reading comprehension
xMRC is a relatively new task of natural language understanding. Cui et al. (2019) proposed the task, in order to improve non-English MRC performance by introducing English resources. There are some representative datasets in this area, such as XQA Liu et al. (2019), BiPaR Jing et al. (2019), MLQA Lewis et al. (2020), and XQuAD Artetxe et al. (2020). Ushio et al. (2023) also proposes a pipeline for multilingual QA generation. However, previous work on the xMRC task mainly focuses on enhancing the performance of task-specific models using techniques such as data augmentation Bornea et al. (2021); Xiang et al. (2024), knowledge injection Duan et al. (2021), constractive learning Chen et al. (2022) and knowledge transfer Cao et al. (2023); Xu et al. (2023). In this paper, we study the xMRC of multilingual LLMs under the larger topic of cross-lingual alignment.
@@ 7 Conclusion
This paper investigates the performance and mechanism of cross-lingual context retrieval of LLMs within the xMRC scenario. For evaluation, we demonstrate the strong xMRC abilities of post-trained models, and study their bottlenecks and oracle performances. For mechanism, we verify a two-phased hypothesis of the xMRC process, identifying the effect of post-training, and finding a possible explanation to the language gap for larger post-trained models. We hope our research will inspire future study to foster the cross-lingual alignment of LLMs in a broader scope.
@@
While this study provides insights of the cross-lingual context retrieval abilities of LLMs, there are also some limitations.
First, the scope of our empirical evaluation is constrained by available resources and time. This necessarily limits the breadth of our testing, preventing us from exhaustively covering the rapidly expanding landscape of LLMs. Besides, while we test across 12 diverse languages, a more comprehensive analysis would ideally include an even wider range of languages, as well as grouping them according to their levels of resource, in order to improve the generalizability of our findings across linguistic diversity.
Second, although we identify a two-phased feature of xMRC and confirm its correlation with pre-training and post-training, the precise factors within these training processes that drive this outcome remain unclear. Future work could delve deeper into the data and strategies of these training stages to locate factors contributing to the emergence and strength of the phasing.
Finally, within the two major phases we discover, we observe hints of more fine-grained changes in model behavior, particularly in the hidden state similarity curves. These preliminary observations suggest the potential for a more nuanced understanding of the xMRC process. Future studies could further investigate these finer-grained dynamics within each phase to gain a more detailed and complete picture of how LLMs achieve cross-lingual context retrieval.
@@
Our use of these existing artifacts, including both language models and datasets, is aligned with their intended use within research contexts. Specifically, derivatives of data accessed for research purposes, such as model outputs and analysis results, are used solely within the bounds of academic inquiry and are not disseminated or utilized outside of these research contexts, in accordance with responsible data handling practices.
@@ Acknowledgments
We would like to thank the anonymous reviewers for their insightful comments. Shujian Huang is the corresponding author. This work is supported by National Science Foundation of China (No. 62376116, 62176120), research project of Nanjing University-China Mobile Joint Institute (NJ20250038), the Fundamental Research Funds for the Central Universities (No. 2024300507, 2025300390).
Appendix A Additional experiment information
A.1 Example of unnecessary source text in outputs
We do not use exact matching (EM) as the main performance metric because it can be easily affected by unnecessary source text in the model outputs. Here is an example:
Context: “The Panthers defense gave up just 308 points, ranking sixth in the league, while also leading the NFL in interceptions with 24 and boasting four Pro Bowl selections. Pro Bowl defensive tackle Kawann Short led the team in sacks with 11, while also forcing three fumbles and recovering two. Fellow lineman Mario Addison added 6½ sacks. The Panthers line also featured veteran defensive end Jared Allen, a 5-time pro bowler who was the NFL’s active career sack leader with 136, along with defensive end Kony Ealy, who had 5 sacks in just 9 starts. Behind them, two of the Panthers three starting linebackers were also selected to play in the Pro Bowl: Thomas Davis and Luke Kuechly. Davis compiled 5½ sacks, four forced fumbles, and four interceptions, while Kuechly led the team in tackles (118) forced two fumbles, and intercepted four passes of his own. Carolina’s secondary featured Pro Bowl safety Kurt Coleman, who led the team with a career high seven interceptions, while also racking up 88 tackles and Pro Bowl cornerback Josh Norman, who developed into a shutdown corner during the season and had four interceptions, two of which were returned for touchdowns.” Question: How many Panthers defense players were selected for the Pro Bowl? Reference: four Model Answer: four Pro Bowl selections.
The reference is “four”, which is consistent with the model output, but the “Pro Bowl selections.” in the model output is unnecessary and will cause the EM metric to give a 0 result.
A.2 Accuracy of language detection tool
We use Lingua with its high accuracy mode in this work to detect the language of model outputs. Table 3 shows its reported accuracies on the tested languages, which is satisfactory to serve in our experiments.
Code Language
Average Acc
(high accuracy mode)
en English 81
de German 89
es Spanish 70
vi Vietnamese 91
zh Chinese 100
hi Hindi 73
ar Arabic 98
el Greek 100
ro Romanian 87
ru Russian 90
th Thai 99
tr Turkish 94
Table 3: Detection accuracies of the tested language with Lingua in its high-accuracy mode, adapted from their GitHub page (https://github.com/pemistahl/lingua).
A.3 Prompt used in evaluation and error type detection
We use the default system prompt and chat templates assigned in the tokenizer.config files of the model repositories. Then, we apply these configurations to two prompt formats which we call v1 and v2.
The v1 prompt format is:
The v2 prompt format is:
The prompt for error type detection is:
<|im_start|>
user
Appendix B More evaluation results
B.1 Full list of models evaluated